Rachel will rent a car for the weekend. She can choose one of the two plans. The first plan has an initial fee of $48 and costs an additional $0.16 per mile driven. The second plan has an initial fee of $57 and costs an additional $0.14 per mile driven.

For what amount of driving do the two plans cost the same?
What is the cost when the two plans cost the same?

Answer by Theo(13342)

plan 1 cost = 48 + .16 * x
plan 2 cost = 57 + .14 * x
x is the number of miles.
plan 1 cost will be the same as plan 2 cost when:
48 + .16 * x = 57 + .14 * x
subtract .14 * x from both sides of the equation and subtract 48 from both sides of the equation to get:
.16 * x - .14 * x = 57 - 48
combine like terms to get:
.02 * x = 9
divide both sides of the equation by .02 to get:
x = 9 / .02 = 450
the two plans should break even at 450 miles driven.
at 450 miles driven,.....
plan 1 cost = 48 + .16 * 450 = 120.
plan 2 cost = 57 + .14 * 450 = 120.

your solution is:
the plans will cost the same when 450 miles have been driven.
the cost when both plans cost the same is 120 collars.